Gross Domestic Product

Use the data chart to answer the questions below.

Consumption

$500

Government Expenditures

$200

GNP

$800

Gross Private Domestic Investment

$100

Imports

$100

Exports

$50


Questions:

1.    Using the data above, GDP is equal to __________________.

2.    Using the data above, Net Exports is equal to _____________.

3.    Using the data above, PCE (Personal Consumption Expenditure) is equal to _____________.

4.    If the dollar amounts of the items above are the nominal amounts in year 2017, and the quantities of units are identical in years 2010 and 2017, but prices in 2010 were 20% less than 2017, Real GDP in 2017 using 2010 as the base year is equal to _______________.
Answer #1

1.

In the open economy, the GDP can be defined the market value of all goods and services which are produced in the domestic territory of the country in the current financial years. GDP by Expenditure method

GDP=C+I+G+(X-M)

=500+100+200+50-100

=750

2.

Net export=X-M

=50-100

=-50

3.

Personal consumption expenditure=$500

4.

Real GDP = (nominal GDP / GDP deflator)*100

GDP deflator= (Nominal GDP/ real GDP)*100

2017 Nominal GDP=750

In 2010, the price were 20% less than the 2017 price level.

Hence real GDP in 2017= Nominal GDP in 2017* 0.80

=750*0.80

=$600